BD - diagonal 1 AB = 1.BC = 3,300 DC 12. In quadrilateral ABCD, RC I AR and determine DA
Answer #1

1 д ч Now Ac= Ag + 3с- де? - y? 132- Ас? = 2s AC = 2х = 5 ДС = 5 . Now in DACO -А? - 421 DC2- р: s2 1122 ДО 2 + 1 ч лю = 109